# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks but remain fundamentally static, unable to adapt their internal parameters to novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this static nature has become a critical bottleneck, necessitating agents that can adaptively reason, act, and evolve in real time. This paradigm shift —from scaling static models to developing self-evolving agents — has sparked growing interest in architectures and methods enabling continual learning and adaptation from data, interactions, and experiences. This survey provides the first systematic and comprehensive review of self-evolving agents, organizing the field around three foundational dimensions — what to evolve, when to evolve, and how to evolve. We examine evolutionary mechanisms across agent components (e.g., models, memory, tools, architecture), categorize adaptation methods by stages (e.g., intra-test-time, inter-test-time), and analyze the algorithmic and architectural designs that guide evolutionary adaptation (e.g., scalar rewards, textual feedback, single-agent and multi-agent systems). Additionally, we analyze evaluation metrics and benchmarks tailored for self-evolving agents, highlight applications in domains such as coding, education, and healthcare, and identify critical challenges and research directions in safety, scalability, and co-evolutionary dynamics. By providing a structured framework for understanding and designing self-evolving agents, this survey establishes a roadmap for advancing adaptive, robust, and versatile agentic systems in both research and real-world deployments, ultimately shedding lights to pave the way for the realization of Artificial Super Intelligence (ASI), where agents evolve autonomously, performing at or beyond human-level intelligence across a wide array of tasks.
Figure 1: A conceptual trajectory illustrating the progression from large language models (LLMs) to foundation agents, advancing to self-evolving agents—our focus, and ultimately toward hypothetical Artificial Super Intelligence (ASI). Along this path, intelligence and adaptivity increase, marking a shift toward more autonomous and agentic AI systems.
@@ 1 Introduction
"It is not the most intellectual of the species that survives; it is not the strongest that survives; but the species that survives is the one that is able best to adapt and adjust to the changing environment in which it finds itself."                          — Charles Darwin [^1] This quote is widely attributed to Charles Darwin, but it does not appear verbatim in his writings. The phrasing is believed to originate from Professor Leon C. Megginson, who paraphrased Darwin’s ideas. Despite its frequent misattribution, the quote effectively captures the essence of Darwinian evolution and has since been popularized in both scientific and managerial literature.
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Yet, they remain fundamentally static luo2025largelanguagemodelagent , unable to adapt their internal parameters when encountering novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this limitation becomes a critical bottleneck. In such settings, conventional knowledge retrieval mechanisms prove inadequate, giving rise to agents capable of dynamically adapting their perception, reasoning, and actions in real time. This emerging need for dynamic, continual adaptation signals a conceptual shift in artificial intelligence: from scaling up static models to developing self-evolving agents, such agents are capable to continuously learn from new data, interactions, and experiences in real-time, leading to systems that are more robust, versatile, and capable of tackling complex, dynamic real-world problems (tool_tut, ). This shift is currently driving us toward a promising and transformative path to Artificial Super Intelligence (ASI), where the agents not only can learn and evolve from experience with an unpredictable speed but also perform at or above human-level intelligence across a wide array of tasks (wang2025theoryagentstoolusedecisionmakers, ).
@@
Unlike static LLMs, which remain constrained by their inability to adapt to novel and evolving contexts, self-evolving agents are designed to overcome these limitations by continuously learning from real-world feedback. This progression reshapes our understanding of agents. Self-evolving agents, as a core concept, will be the precursors to ASI, acting as intermediaries that pave the way for the ultimate evolution of intelligence, as shown in Figure 1. Recent research initiatives have increasingly focused on developing adaptive agent architectures capable of continually learning and adapting from experience, such as recent advancements in agent frameworks (yin2025godelagentselfreferentialagent, ), prompting strategies fernando2023promptbreeder , and different optimization ways to evolve. Notwithstanding these advances, existing surveys predominantly address agent evolution as a subsidiary component within comprehensive agent taxonomies. Previous surveys primarily provide systematic overviews of general agent development, while offering limited coverage of self-evolving mechanisms across constrained scenarios in self-evolving agents (luo2025largelanguagemodelagent, ; liu2025advanceschallengesfoundationagents, ). For example, Luo et al. (luo2025largelanguagemodelagent, ) discuss several ways to evolve, such as self-learning and multi-agent co-evolution, while Liu et al. (liu2025advanceschallengesfoundationagents, ) explicitly introduce the evolution in terms of different components of agents, such as tools and prompts. Moreover, some studies focus specifically on the evolution of language models themselves tao2024survey-self-evo-llm , rather than on the broader concept of agents. Yet, there is no systematic survey devoted to a dedicated, comprehensive investigation of self-evolving agents as a first-class research paradigm. This gap has left fundamental questions underexplored: What aspects of an agent should evolve? When should adaptation occur? And how should that evolution be implemented in practice?
@@
To the best of our knowledge, this is the first systematic and comprehensive survey focusing on self-evolving agents, offering a clear roadmap for both theoretical inquiry and practical deployment. We organize our analysis around three foundational questions — what, when, and how to evolve — and provide a structured framework for understanding each. Specifically, we systematically examine individual agent components, including the model, memory, tools and corresponding workflow, investigating their distinct evolutionary mechanisms (what to evolve of agent in Section 3); then we divide existing evolving methods according to different temporal stages with different learning paradigms such as supervised fine-tuning, reinforcement learning and inference-time evolving (when to evolve in Section 4). We finally summarize different signals to guide the evolution of agents, such as textual feedback or scalar rewards, and also different architectures of agents to evolve, such as single-agent and multi-agent evolution (how to evolve in Section 5). Furthermore, we review certain evaluation metrics and benchmarks to track existing advancements of self-evolving agents, emphasizing the importance of co-evolution between evaluation and agents (Section 6). We also examine emerging applications in domains such as coding, education, and healthcare, where continual adaptation and evolution are essential (Section 7). Finally, we identify persistent challenges and outline promising research directions to guide the development of self-evolving agents (Section 8). Through this systematic decomposition of self-evolutionary processes across orthogonal dimensions, we provide a structured and practical framework enabling researchers to systematically analyze, compare, and design more robust and adaptive agentic systems. To sum up, our key contributions are as follows:
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=0.1em, edge+=darkgray, line width=0.8pt, s sep=2pt, inner xsep=2pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=3em,font=, where level=2text width=5em,font=, where level=3text width=3em,font=, where level=4text width=14em,font=, [ Self-evolving Agent, draw=gray, color=gray!100, fill=gray!15, thick, text=black, ver, [What to evolve, cyan!100, fill=cyan!8,thick, text=black, [Model [ Policy [ SCA zhou2025selfchallenginglanguagemodelagents , Self-Rewarding Self-Improving simonds2025selfrewardingselfimproving , SELF lu2023self , SCoRe kumar2024training , PAG jiang2025pag , TextGrad yellamraju2024textgrad , AutoRule wang2025autorule , SRLM wang2025selfreasoninglanguagemodelsunfold ]] [ Lesson [ SCA zhou2025selfchallenginglanguagemodelagents , AgentGen hu2024agentgen , Reflexion shinn2023reflexion , AdaPlanner sun2023adaplanner , SICA robeyns2025selfimprovingcodingagent , Self-Refine madaan2023selfrefine , Learn-by-interact su2025learnbyinteractdatacentricframeworkselfadaptive , RAGEN wang2025ragen , DYSTIL wang2025dystildynamicstrategyinduction ]]] [Context [ Memory [ SAGE liang2024sage , Mem0 chhikara2025mem0 , MemInsight salama2025meminsight , REMEMBER zhang2024large , Expel zhao2024expel , Agent Workflow Memory wang2024agent , Richelieu diplomacy agent guan2024richelieu , ICE qian2024investigate ]][ Prompt [ APE zhou2022large , ORPO yang2023large , ProTeGi pryzant2023automatic , PromptAgent wang2023promptagent , REVOLVE zhang2024revolve , PromptBreeder fernando2023promptbreeder , DSPy khattab2023dspy , Trace wang2023trace , TextGrad yellamraju2024textgrad , SPO xiang2025self , LLM-AutoDiff yin2025llm , EvoAgent yuan2024evoagent ]]] [Tool[ Creation [ Voyager wang2023voyager , Alita qiu2025alita , ATLASS Haque2025ATLASS , CREATOR qian2023creator , SkillWeaver zheng2025skillweaver , CRAFT yuan2024craft ]][ Mastery [ LearnAct zhao2024learnact , DRAFT fromexplorationtomastery , ToolLLM qin2023toolllm , Toolformer schick2023toolformer , Gorilla patil2024gorilla ]] [ Selection [ ToolGen wang2025toolgen , AgentSquare shang2025agentsquare , Darwin Godel Machine zhang2025darwin , COLT qu2024colt , TOOLRET shi2025toolret , ToolRerank zheng2024toolrerank , PTR gao2024ptr , SSO Nottingham2024sso ]]] [Architecture[ Single-Agent [ AgentSquare shang2025agentsquare , Darwin Godel Machine zhang2025darwin , Gödel Agent  yin2025godelagentselfreferentialagent , AlphaEvolve novikov2025alphaevolve , TextGrad yellamraju2024textgrad , EvoFlow zhang2025evoflow , MASS zhou2025maasprompt ]][ Multi-Agent [ AFlow zhang2024aflow , ADAS hu2024automated , AutoFlow li2024autoflow , GPTSwarm zhuge2024gptswarm , ScoreFlow wang2025scoreflow , FlowReasoner gao2025flowreasoner , ReMA wan2025remalearningmetathinkllms , GIGPO feng2025groupingrouppolicyoptimizationllm ] ] ] ] [ When to evolve, color=green!80!black!50!, fill=green!5, thick, text=black, [ Intra-test-time self-evolution [ ICL [ Reflexion shinn2023reflexion , SELF madaan2023self_refine , AdaPlanner sun2023adaplanner , TrustAgent hua2024trustagent ] ] [ SFT [ Self-Adaptive LM zweiger2025self , TTT-NN hardt2023test , SIFT hubotter2024efficiently ] ] [ RL [ LADDER simonds2025ladder , Ttrl zuo2025ttrl ] ] ] [ Inter-test-time self-evolution [ SFT [ SELF lu2023self , STaR zelikman2022star , Quiet-STaR zelikman2024quiet , SiriuS zhao2025sirius , Chen et al. chen2025sft ] ] [ RL [ RAGEN wang2025ragen , Learning-Like-Humans zhang2025learning , WebRL qi2024webrl , DigiRL bai2024digirl ] ] ] ] [ How to evolve, color=yellow!80!brown, fill=yellow!20!white, thick, text=black [ Reward-based Self-Evolution [Textual Feedback [ Reflexion shinn2023reflexion , AdaPlanner sun2023adaplanner , AgentS2 agashe2025agents2 , SELF lu2023self , Self-Refine madaan2023self_refine , SCoRe kumar2024training , PAG jiang2025pag , TextGrad yellamraju2024textgrad ] ] [Internal Rewards [ CISC taubenfeld2025cisc , Self-Ensemble xu2025selfensemblemitigatingconfidencedistortion , SRSI simonds2025selfrewardingselfimproving , Self-Certainty kang2025scalablebestofnselectionlarge , Self-Rewarding Language Models yuan2025selfrewardinglanguagemodels ] ] [External Rewards [ Self-Train LM shafayat2025largereasoningmodelsselftrain , MM-UPT wei2025unsupervisedposttrainingmultimodalllm , CoVo zhang2025consistentpathsleadtruth , SWE-Dev du2025swedevevaluatingtrainingautonomous , SICA robeyns2025self , Feedback Friction jiang2025feedbackfriction , USEagent applis2025useagent , DYSTIL wang2025dystildynamicstrategyinduction , OTCPO wang2025otcpo , AutoRule wang2025autorule , EGSR zhang2025learning , LADDER simonds2025ladder , RAGEN wang2025ragen , SPIRAL liu2025spiral ] ] [Implicit Rewards [ Reward Is Enough song2025reward , Endogenous reward li2025generalist ] ] ] [ Imitation & Demonstration Learning [Self-Generated [ STaR zelikman2022star , V-STaR hosseini2024v , AdaSTaR koh2025adastar , STIC deng2024enhancing , GENIXER zhao2024genixer ] ] [Others [ SiriuS zhao2025sirius , SOFT tang2025bridging , RISE qu2024recursive , IoE li2024confidence ] ] ] [ Population-based & Evolutionary Methods [Single Agent [ DGM zhang2025darwin , GENOME zhang2025nature , SPIN chen2024self , SPC chen2025spc , STL mendes2025language ] ] [Multi-Agent [ EvoMAC hu2024self , Puppeteer dang2025multi , MDTeamGPT chen2025self , MedAgentSim almansoori2025self ] ] ] ] [ Where to Evolve, color=violet!60, fill=violet!15, thick, text=black, [ General Domain [ Memory [ Mobile-Agent-E wang2025mobile , MobileSteward liu2025mobilesteward ] ] [ Curriculum Learning [ WebRL qi2024webrl , Voyager wang2023voyager ] ] [ Model-Agent Co-Evolution [ WebEvolver fang2025webevolver , UI-Genie xiao2025ui , Absolute-Zero zhao2025absolute ] ] ] [ Specialized Domain [ Coding, GUI Finalicial… [ EvoMac hu2024self , SICA robeyns2025self , AgentCoder dang2025multi , QuantAgent wang2024quantagent ] ] [ Others [ Arxiv-copilot lin2024arxivcopilot , Voyager wang2023voyager ] ] ] ] ]
@@
Figure 3: A comprehensive overview of self-evolving agents across key dimensions: What to evolve: covering four major categories—Model, Context, Tool, and Architecture; When to evolve: differentiating intra-test-time and inter-test-time self-evolution, via in-context learning (ICL), supervised fine-tuning (SFT), or reinforcement learning (RL); How to evolve: centered on three main paradigms—reward-based, imitation and demonstration, and population-based methods. These are complemented by cross-cutting dimensions. Where to evolve: ranging from general-purpose domains to specific domains; Evaluation: focusing on goals (e.g., adaptivity, safety, generalization) and evaluation paradigms (static, short-horizon, or long-horizon).
@@ 2.1 Definitions
[…]
@@
[…]
@@
We define a (multi-)agent system as $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$. The architecture $\Gamma$ determines the control flow of the agent system or collaborative structures between multiple agents. It is typically represented as a sequence of nodes $(N_{1},N_{2},...)$ organized by graph or code structures. Each node $N_{i}$ consists of the following components:
@@
For a given task $\mathcal{T}=(E,g)$, represented by an environment $E$ and a corresponding goal $g\in\mathcal{G}$, the agent system follows the topology $\Gamma$ to generate a trajectory $\tau=(o_{0},a_{0},o_{1},a_{1},...)$, and receives a feedback $r$ either from the external environment or from internal signals (e.g., self-confidence or feedback from an evaluator).

### v4

# A Survey of Self-Evolving Agents What, When, How, and Where to Evolve on the Path to Artificial Super Intelligence

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks but remain fundamentally static, unable to adapt their internal parameters to novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this static nature has become a critical bottleneck, necessitating agents that can adaptively reason, act, and evolve in real time. This paradigm shift —from scaling static models to developing self-evolving agents — has sparked growing interest in architectures and methods enabling continual learning and adaptation from data, interactions, and experiences. This survey provides the first systematic and comprehensive review of self-evolving agents, organizing the field around three foundational dimensions — what to evolve, when to evolve, and how to evolve. We examine evolutionary mechanisms across agent components (e.g., models, memory, tools, architecture), categorize adaptation methods by stages (e.g., intra-test-time, inter-test-time), and analyze the algorithmic and architectural designs that guide evolutionary adaptation (e.g., scalar rewards, textual feedback, single-agent and multi-agent systems). Additionally, we analyze evaluation metrics and benchmarks tailored for self-evolving agents, highlight applications in domains such as coding, education, and healthcare, and identify critical challenges and research directions in safety, scalability, and co-evolutionary dynamics. By providing a structured framework for understanding and designing self-evolving agents, this survey establishes a roadmap for advancing more adaptive, capable, robust, and versatile agentic systems in both research and real-world deployments, and ultimately sheds light on the realization of Artificial Super Intelligence (ASI) where agents evolve autonomously and perform beyond human-level intelligence across a wide array of tasks.
Figure 1: A conceptual trajectory illustrating the progression from large language models (LLMs) to foundation agents, and then to self-evolving agents—our focus, and ultimately toward the hypothetical Artificial Super Intelligence (ASI). Along this path, intelligence and adaptivity increase, marking a shift toward more autonomous and agentic AI systems. The future directions beyond self-evolving agents remain open and subject to ongoing exploration.
@@ 1 Introduction
"It is not the most intellectual of the species that survives; it is not the strongest that survives; but the species that survives is the one that is able best to adapt and adjust to the changing environment in which it finds itself."                 ― Charles Darwin [^1] This quote is widely attributed to Charles Darwin, but it does not appear verbatim in his writings. The phrasing is believed to originate from Professor Leon C. Megginson, who paraphrased Darwin’s ideas. Despite its frequent misattribution, the quote effectively captures the essence of Darwinian evolution and has since been popularized in both scientific and managerial literature.
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Yet, they remain fundamentally static (Luo et al., 2025a), unable to adapt their internal parameters when encountering novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this limitation becomes a critical bottleneck. In such settings, conventional knowledge retrieval mechanisms prove inadequate, giving rise to agents capable of dynamically adapting their perception, reasoning, and actions in real time. This emerging need for dynamic, continual adaptation signals a conceptual shift in artificial intelligence: from scaling up static models to developing self-evolving agents. While established techniques like Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) provide the mechanisms for improvement, we define self-evolution not merely by the algorithms used, but by the locus of autonomy. Unlike traditional pipelines where human engineers curate data and schedule updates, a self-evolving agent is capable of continuously learning from new data, interactions, and experiences in real-time, leading to systems that are more robust, versatile, and capable of tackling complex, dynamic real-world problems (Wang et al., 2024a). This shift is currently driving us toward a promising and transformative path to Artificial Super Intelligence (ASI), where the agents not only can learn and evolve from experience with an unpredictable speed but also perform at or above human-level intelligence across a wide array of tasks (Wang et al., 2025f).
@@
Unlike static LLMs, which remain constrained by their inability to adapt to novel and evolving contexts, self-evolving agents are designed to overcome these limitations by continuously learning from real-world feedback. This progression reshapes our understanding of agents. Self-evolving agents, as a core concept, represent a significant step forward in the evolution of intelligent systems, acting as intermediaries that pave the way for more adaptive and autonomous AI, as shown in Figure 1. Recent research initiatives have increasingly focused on developing adaptive agent architectures capable of continually learning and adapting from experience, such as recent advancements in agent frameworks (Yin et al., 2025), prompting strategies (Fernando et al., 2023), and different optimization ways to evolve. Notwithstanding these advances, existing surveys predominantly address agent evolution as a subsidiary component within comprehensive agent taxonomies. Previous surveys primarily provide systematic overviews of general agent development, while offering limited coverage of self-evolving mechanisms across constrained scenarios in self-evolving agents (Luo et al., 2025a; Liu et al., 2025a). For example,  Luo et al. (2025a) discuss several ways to evolve, such as self-learning and multi-agent co-evolution, while Liu et al. (2025a) explicitly introduce the evolution in terms of different components of agents, such as tools and prompts. Moreover, some studies focus specifically on the evolution of language models themselves (Tao et al., 2024), rather than on the broader concept of agents. However, these works address isolated components rather than the holistic agent system. Therefore, there is no systematic survey devoted to a dedicated, comprehensive investigation of self-evolving agents as a first-class research paradigm. This gap has left fundamental questions underexplored: What aspects of an agent should evolve? When should adaptation occur? And how should that evolution be implemented in practice?
@@
To the best of our knowledge, this is the first systematic and comprehensive survey focusing on self-evolving agents, offering a clear roadmap for both theoretical inquiry and practical deployment. However, given that this represents a rapidly forming research area where conceptual boundaries are still being actively negotiated within the community, we frame this survey as a guiding synthesis rather than a review of a fully established paradigm. Instead of enforcing rigid boundaries, we aim to structure the heterogeneous mechanisms emerging in the community into a coherent framework. We organize our analysis around three foundational questions — what, when, and how to evolve — and provide a structured framework for understanding each. Specifically, we systematically examine individual agent components, including the model, memory, tools and corresponding workflow, investigating their distinct evolutionary mechanisms (what to evolve of agent in Section 3); then we divide existing evolving methods according to different temporal stages with different learning paradigms such as supervised fine-tuning, reinforcement learning and inference-time evolving (when to evolve in Section 4). We finally summarize different signals to guide the evolution of agents, such as textual feedback or scalar rewards, and also different architectures of agents to evolve, such as single-agent and multi-agent evolution (how to evolve in Section 5). Furthermore, we review certain evaluation metrics and benchmarks to track existing advancements of self-evolving agents, emphasizing the importance of co-evolution between evaluation and agents (Section 6). We also examine emerging applications in domains such as coding, education, and healthcare, where continual adaptation and evolution are essential (Section 7). Finally, we identify persistent challenges and outline promising research directions to guide the development of self-evolving agents (Section 8). Through this systematic decomposition of self-evolutionary processes across orthogonal dimensions, we provide a structured and practical framework enabling researchers to systematically analyze, compare, and design more robust and adaptive agentic systems. To sum up, our key contributions are as follows:
@@
Figure 3: A comprehensive overview of self-evolving agents across key dimensions. From left to right and top to bottom, the figure mirrors the organization of Sections 37. What to evolve (Sec. 3) decomposes agent components: model, context, tools, and architecture, showing where evolution operates. When to evolve (Sec. 4) distinguishes intra-test-time versus inter-test-time self-evolution, corresponding to ICL, SFT, and RL paradigms. How to evolve (Sec. 5) summarizes methodological families—reward-based, imitation & demonstration, and population-based—together with cross-cutting dimensions such as online/offline, on/off-policy, and reward granularity. Where to evolve (Sec. 6) contrasts general-purpose and domain-specific deployments (e.g., coding, GUI, finance, medical, education). Evaluation (Sec. 7) outlines goals and metrics—adaptivity, generalization, efficiency, safety—and corresponding evaluation paradigms (static, short-horizon, long-horizon). Overall, the taxonomy maps the survey’s reasoning flow: defining what, when, and how to evolve establishes the foundation for evaluating and advancing self-evolving agents.
@@ 2.1 Definitions
  • $T$ is the state transition probability function which takes a state-action pair $(s,a)$ and outputs the probability distribution $T(s^{\prime}|s,a)$ of the next state.
@@
  • $O$ is the observation probability function which takes a state-action pair $(s,a)$ and outputs the probability distribution $O(o^{\prime}|s,a)$ of the next observation for the agent.
@@
[…]
@@
[…]
@@
Operational definition of self-evolving agents
To provide a conceptual boundary, we introduce an operational definition of self-evolving agents. A self-evolving agent is the agent that modifies its internal parameters, contextual state, toolset, or architectural topology based on its own trajectories or feedback signals, with the explicit objective of improving future performance.
This definition entails three inclusion criteria: (i) updates must be experience-dependent, driven by trajectories, self-generated data, or environment feedback, specifically targeting the agent’s policy limitations or capability boundaries rather than generic data synthesis; (ii) updates must produce a persistent, policy-changing effect rather than a transient instruction-following behavior; (iii) the system must possess mechanisms for autonomous exploration or self-initiated learning, even if it also leverages pre-collected data. For clarity, we use "passive" to denote learning triggered exclusively by externally provided data or schedules, and "active" to denote self-initiated exploration, reflection, or structural modification (i.e., using self-reflection to collect data), explicitly excluding static pipelines (e.g., standard distillation) where data generation is agnostic to the agent’s interaction history.
As this field is rapidly forming, fully autonomous self-evolution without human intervention represents an aspirational goal rather than the current norm. In this survey, we do not impose a rigid exclusion threshold that would disregard early-stage developments. Instead, we analyze the mechanisms contributing to the self-evolving paradigm ranging from proto-evolution (e.g., iterative bootstrapping or feedback-driven prompting) to strong self-evolution (fully autonomous diagnosis and reconfiguration), allowing us to provide a comprehensive view of how diverse methods contribute to the "What, When, and How" of the paradigm’s progression toward full autonomy.